OpenStudy (anonymous):

In a certain dice game, three fair dice are rolled and the total on all three dice is counted. Let the random variable x be the total count on all three dice. You pay one dollar to roll the dice. If you roll a 16 or greater, you will win $10. If you roll a 6 or lower, you win $5. If you roll anything else, you lose your money. What is the expected value of this game?

OpenStudy (anonymous):

Well we know \(X\) always gives some value on the interval \([3\ ..\,36]\), with \(34\) different sums that are all equally likely. We know that for \(X\ge16\), our payout \(P\) will be \(10\) (this happens with probability \(21/34\) since there are \(21\) possible numbers that work here). For \(X\le6\), we have \(P=5\) (this happens with probability \(4/34=2/17\) since there are \(4\) possible numbers that work here). We also have the event that \(6\lt X\lt 16\), we have \(P=0\) (probability \(9/34\)).

OpenStudy (anonymous):

Our *expected value* of \(P\) is just an average of its possible values, each weighed by their probability, so:$$E(P)=\frac{21}{34}\times10+\frac2{17}\times5+\frac9{34}\times0=\frac{115}{17}$$

OpenStudy (anonymous):

weighTed

OpenStudy (anonymous):

Thank you!!!!